A 0.22 kg piece of silly putty is thrown at a 1.2 kg wooden panel that is initially at the rest. The velocity of the silly putty before the collision was 18.5 m/s. The silly putty sticks to the wooden panel and they move as one after the collision. What is the velocity of the combined objects after the collision.
1
Expert's answer
2020-11-05T10:45:36-0500

According to momentum conservation, the initial momentum equals the final:


"p_i=p_f,\\
m_sv_s+m_wv_w=(m_s+m_w)u,\\\space\\
u=\frac{m_sv_s+m_wv_w}{m_s+m_w}=\frac{0.22\cdot18.5+1.2\cdot0}{0.22+1.2}=2.9\text{ m/s}."